OpenStudy (anonymous):

2sin^2 + 5cosx=4 , solve for x

OpenStudy (mertsj):

Replace sin^2x with 1-cos^2x

OpenStudy (dumbcow):

^^ after this solve it as a quadratic equation, you could say u = cos x then 2u^2 -5u +2 = 0

OpenStudy (anonymous):

isnt it 2x^2 -5x-3=0

OpenStudy (loser66):

Mertsj suggests replace sin^2 = 1 -cos ^2 . So you have 2 sin^2 = 2 -2cos^2 . Replace to your equation you have 2 -2cos^2 + 5cos =4 ---> -2cos^2 +5cos -2 =0 or cos^2 -5cos +2 =0. let u = cos . you have 2u^2 -5u +2=0 don't know how you get 3. Recheck.

OpenStudy (mertsj):